OpenStudy (ars22):

(a + b) - (a + b)^2 ---------------- (a + b) (1 - a - b simplify

OpenStudy (callisto):

\[\frac{(a + b) - (a + b)^2}{(a + b) (1 - a - b)}\]\[=\frac{(a + b)[1- (a + b)]}{(a + b) (1 - a - b)}\]\[=\frac{(a + b)(1- a - b)}{(a + b) (1 - a - b)}\]Can you simplify it by cancelling the common factors there?

OpenStudy (ars22):

thanks

OpenStudy (callisto):

welcome!~